HIGH-RESOLUTION FOURIER-TRANSFORM INTRA-CAVITY LASER ABSORPTION SPECTROSCOPY: APPLICATION TO ^{12}C_{2}H_{2} NEAR 12300 cm^{−1}

\underline{SHUI-MING HU}, ZHI-YONG WU, YUN DING, AN-WEN LIU, Laboratory of Bond Selective Chemistry, University of Science and Technology of China, Hefei, 230026, China; ALAIN CAMPARGUE, Laboratoire de Spectrométrie Physique (UMR5588), Université Joseph Fourier de Grenoble, BP 87, 38402 Saint Martin d’Héres, France.

The capabilities of ICLAS associated with a high resolution Fourier Transform spectrometer (FT-ICLAS) are investigated with a Ti : Sapphire laser. Weak absorption lines of atmospheric water were used to test the accuracy of absolute intensity measurements by FT-ICLAS leading to an excellent agreement (a few %) with the intensity values obtained by conventional FTS. The performances in terms of spectral resolution (0.028 cm^{−1}) and sensitivity (α_{min} ≈ 2 × 10^{−9} cm^{−1}) are illustrated by the spectroscopic study of the overtone spectrum of^{12}C_{2}H_{2} between 12250 and 12400 cm^{−1} which allowed for a significant improvement of recent CRDS measurements. Among the three Π − Σ bands rotationally analyzed, one is newly observed. The absolute intensity values of the bands are given.